### v1

@@ Methods
To reduce the number of low-energy constants involved in the $\Xi NN$ potentials, we employ the decuplet saturation approximation, by which only two of them remain unconstrained.
@@ Conclusions
High-momentum scattering of the deuteron off $\Xi^{-}$ is expected to be a promising probe of the $\Xi NN$ 3BF, offering an alternative to femtoscopic analyses of the correlation function.
@@ I Introduction
Three-baryon forces (3BFs) play a crucial role in understanding the properties of hypernuclei. For instance, the impact of $\Lambda NN$ 3BF on hyperneutron matter was reported in Ref. [1], where the $\Lambda NN$ 3BF was modeled as a two-pion exchange (TPE) plus phenomenological repulsive term. More recently, lattice simulations of hyperneutron matter revealed the effect of $\Lambda NN$ and $\Lambda\Lambda N$ 3BFs, which are defined under the pionless effective field theory (EFT) [2, 3].
@@
This paper is organized as follows. In Sect. II, we formulate the general expression of the $\Xi NN$ 3BF as potentials in momentum space. We also give formulae necessary for implementing the potentials in the calculations of the correlation function. In Sect. III, we demonstrate how the $\Xi NN$ 3BF affects the correlation function. Section IV is devoted to the summary and perspectives of this work.
@@ II.1.2 Two-pion exchange term
$v_{\mathrm{TME}}$ $=-\dfrac{1}{4f_{0}^{2}}\dfrac{\left(\bm{\sigma}_{i}\cdot\bm{q}_{li}\right)\left(\bm{\sigma}_{k}\cdot\bm{q}_{nk}\right)}{\left(q_{li}^{2}+m_{\phi_{1}}^{2}\right)\left(q_{nk}^{2}+m_{\phi_{2}}^{2}\right)}\left[N_{1}^{\prime}+N_{2}^{\prime}\bm{q}_{li}\cdot\bm{q}_{nk}+N_{3}^{\prime}i\bm{\sigma}_{j}\cdot(\bm{q}_{li}\times\bm{q}_{nk})\right],$ (2)
@@
with the exchange operator $\mathcal{P}_{ij}=\mathcal{P}^{(q)}_{ij}\mathcal{P}^{(\sigma)}_{ij}\mathcal{P}^{(\tau)}_{ij}$}, where $\mathcal{P}^{(q)}_{ij}$, $\mathcal{P}^{(\sigma)}_{ij}$, and $\mathcal{P}^{(\tau)}_{ij}$ respectively exchange the parity, spin, and isospin of the subsystem formed by baryons $i$ and $j$. The pion decay constant with the finite pion mass is expressed by $f_{\pi}$, for which we employ $f_{\pi}=92.4$ MeV, and $m_{\pi}=138.04$ MeV is the average pion mass. The axial vector coupling constants $g_{A}$ and $g_{B}$ are expressed as $g_{A}=D+F$ and $g_{B}=-D+F$ with $D\approx 0.8$ and $F\approx 0.5$ [6]. The LECs $c_{i}$ and $u_{i}$ are given as a linear combination of other LECs;
@@ II.1.3 One-pion exchange plus contact term
$X_{123}^{456}$ $=\dfrac{g_{B}d}{2f_{\pi}^{4}}\dfrac{\bm{\sigma}_{1}\cdot\bm{q}_{41}}{q_{41}^{2}+m_{\pi}^{2}}\left[(\bm{\tau}_{2}-\bm{\tau}_{3})\cdot\bm{\tau}_{1}\,(\bm{\sigma}_{2}-\bm{\sigma}_{3})\cdot\bm{q}_{41}+(\bm{\tau}_{1}\times\bm{\tau}_{2})\cdot\bm{\tau}_{3}\,(\bm{\sigma}_{2}\times\bm{\sigma}_{3})\cdot\bm{q}_{41}\right],$ (16)
$X_{231}^{564}$ $=\dfrac{g_{A}}{2f_{\pi}^{4}}\dfrac{\bm{\sigma}_{2}\cdot\bm{q}_{52}}{q_{52}^{2}+m_{\pi}^{2}}\left[(e_{1}\bm{\sigma}_{3}+e_{2}\bm{\sigma}_{1})\cdot\bm{q}_{52}\,\bm{\tau}_{2}\cdot\bm{\tau}_{3}+(e_{3}\bm{\sigma}_{3}+e_{4}\bm{\sigma}_{1})\cdot\bm{q}_{52}\,\bm{\tau}_{1}\cdot\bm{\tau}_{2}-e_{5}(\bm{\tau}_{2}\times\bm{\tau}_{3})\cdot\bm{\tau}_{1}\,(\bm{\sigma}_{3}\times\bm{\sigma}_{1})\cdot\bm{q}_{52}\right],$ (17)
$X_{312}^{645}$ $=\dfrac{g_{A}}{2f_{\pi}^{4}}\dfrac{\bm{\sigma}_{3}\cdot\bm{q}_{63}}{q_{63}^{2}+m_{\pi}^{2}}\left[(e_{4}\bm{\sigma}_{1}+e_{3}\bm{\sigma}_{2})\cdot\bm{q}_{63}\,\bm{\tau}_{3}\cdot\bm{\tau}_{1}+(e_{2}\bm{\sigma}_{1}+e_{1}\bm{\sigma}_{2})\cdot\bm{q}_{63}\,\bm{\tau}_{2}\cdot\bm{\tau}_{3}-e_{5}(\bm{\tau}_{3}\times\bm{\tau}_{1})\cdot\bm{\tau}_{2}\,(\bm{\sigma}_{1}\times\bm{\sigma}_{2})\cdot\bm{q}_{63}\right],$ (18)
$X_{312}^{564}$ $=\dfrac{g_{A}}{2f_{\pi}^{4}}\dfrac{\bm{\sigma}_{3}\cdot\bm{q}_{63}}{q_{63}^{2}+m_{\pi}^{2}}\left[\{f_{1}\bm{\sigma}_{1}+f_{2}\bm{\sigma}_{2}+f_{3}i(\bm{\sigma}_{1}\times\bm{\sigma}_{2})\}\cdot\bm{q}_{63}\,\bm{\tau}_{3}\cdot\bm{\tau}_{1}+\{f_{2}\bm{\sigma}_{1}+f_{1}\bm{\sigma}_{2}+f_{3}i(\bm{\sigma}_{1}\times\bm{\sigma}_{2})\}\cdot\bm{q}_{63}\,\bm{\tau}_{2}\cdot\bm{\tau}_{3}\right$.
@@
$X_{231}^{645}$ $=\dfrac{g_{A}}{2f_{\pi}^{4}}\dfrac{\bm{\sigma}_{2}\cdot\bm{q}_{52}}{q_{52}^{2}+m_{\pi}^{2}}\left[\{f_{1}\bm{\sigma}_{3}+f_{2}\bm{\sigma}_{1}-f_{3}i(\bm{\sigma}_{3}\times\bm{\sigma}_{1})\}\cdot\bm{q}_{52}\,\bm{\tau}_{2}\cdot\bm{\tau}_{3}+\{f_{2}\bm{\sigma}_{3}+f_{1}\bm{\sigma}_{1}-f_{3}i(\bm{\sigma}_{3}\times\bm{\sigma}_{1})\}\cdot\bm{q}_{52}\,\bm{\tau}_{1}\cdot\bm{\tau}_{2}\right$.
@@ II.1.4 Contact term
$V_{ijk}^{a}=\begin{dcases}1&(a=1),\\ \bm{\sigma}_{j}\cdot\bm{\sigma}_{k}&(a=2),\\ \bm{\sigma}_{k}\cdot\bm{\sigma}_{i}&(a=3),\\ \bm{\sigma}_{i}\cdot\bm{\sigma}_{j}&(a=4),\\ i\bm{\sigma}_{i}\cdot\left(\bm{\sigma}_{j}\times\bm{\sigma}_{k}\right)&(a=5).\end{dcases}$ (34)
@@
$+\left[F_{1}-F_{2}(\bm{\sigma}_{2}-\bm{\sigma}_{3})\cdot\bm{\sigma}_{1}-F_{1}\bm{\sigma}_{2}\cdot\bm{\sigma}_{3}\right]\bm{\tau}_{1}\cdot\bm{\tau}_{3}+\left[G_{1}-E_{2}(\bm{\sigma}_{2}+\bm{\sigma}_{3})\cdot\bm{\sigma}_{1}-\dfrac{1}{3}E_{1}\bm{\sigma}_{2}\cdot\bm{\sigma}_{3}\right]\bm{\tau}_{2}\cdot\bm{\tau}_{3}.$ (35)
The LECs $E_{i}$, $F_{i}$, and $G_{1}$ can be expressed as a linear combination of the other LECs $C_{i}$, which enter the particle-basis Lagrangian [6];
@@
$W^{(\sigma)}(\bm{r},\bm{R})$ $=\frac{1}{(2\pi)^{6}}\iint d\bm{q}_{1}d\bm{q}_{2}\exp\left[i(\bm{q}_{1}\cdot\bm{r}_{1}+\bm{q}_{2}\cdot\bm{r}_{2})\right]\Braket{\Upsilon_{\sigma m_{\sigma}}\Theta_{1/2,-1/2}|V_{(0)}^{\Xi NN}(\bm{q}_{1},\bm{q}_{2},\bm{q}_{3})|\Upsilon_{\sigma m_{\sigma}}\Theta_{1/2,-1/2}}$
$=\frac{1}{(2\pi)^{6}}\iint d\bm{q}_{2}d\bm{q}_{3}\exp\left[i(\bm{q}_{2}\cdot\bm{r}_{2}+\bm{q}_{3}\cdot\bm{r}_{3})\right]\Braket{\Upsilon_{\sigma m_{\sigma}}\Theta_{1/2,-1/2}|V_{(0)}^{\Xi NN}(\bm{q}_{1},\bm{q}_{2},\bm{q}_{3})|\Upsilon_{\sigma m_{\sigma}}\Theta_{1/2,-1/2}}$
$=\frac{1}{(2\pi)^{6}}\iint d\bm{q}_{3}d\bm{q}_{1}\exp\left[i(\bm{q}_{3}\cdot\bm{r}_{3}+\bm{q}_{1}\cdot\bm{r}_{1})\right]\Braket{\Upsilon_{\sigma m_{\sigma}}\Theta_{1/2,-1/2}|V_{(0)}^{\Xi NN}(\bm{q}_{1},\bm{q}_{2},\bm{q}_{3})|\Upsilon_{\sigma m_{\sigma}}\Theta_{1/2,-1/2}}.$ (46)
@@
$U_{\mathrm{TPE}}^{(u_{1})}(R)$ $=\frac{3g_{A}^{2}u_{1}m_{\pi}^{6}}{32\pi^{2}f_{\pi}^{4}}\sum_{\lambda_{1}=0,1}\sum_{\lambda_{2}=0,1}\sum_{l}(-)^{\lambda}_{1}R^{\lambda_{1}+\lambda_{2}}\widehat{1-\lambda_{1}}\widehat{1-\lambda_{2}}\left[\left(\matrixquantity{3\\ 2\lambda_{1}}\right)\left(\matrixquantity{3\\ 2\lambda_{2}}\right)\right]^{\frac{1}{2}}\left\{\matrixquantity{1-\lambda_{1}&\lambda_{1}&1\\ \lambda_{2}&1-\lambda_{2}&l}\right\}$
@@
$U_{\mathrm{OPE}}^{(\sigma)}(R)$ $=-\dfrac{g_{A}G_{e}^{(\sigma)}m_{\pi}^{3}}{16f_{\pi}^{2}}\int dr\varphi^{2}(r)\int_{-1}^{1}dw\left\{Y(m_{\pi}r_{2})f_{R_{0}}(r_{2})D_{R_{0}}(r_{3})+Y(m_{\pi}r_{3})f_{R_{0}}(r_{3})D_{R_{0}}(r_{2})\right\}$
$\quad-\dfrac{3g_{A}G_{f}^{(\sigma)}m_{\pi}^{3}}{8f_{\pi}^{2}}\int dr\varphi^{2}(r)\int_{-1}^{1}dw\left\{Y(m_{\pi}r_{2})f_{R_{0}}(r_{2})D_{R_{0}}(r)+Y(m_{\pi}r_{3})f_{R_{0}}(r_{3})D_{R_{0}}(r)\right\},$ (49)
@@
Accordingly, the delta functions, which originate from the contact vertices [see Eqs. (105) and (106) in Appendix B], are regularized similarly;
@@
The newly defined LECs $G_{e}^{(\sigma)}$, $G_{f}^{(\sigma)}$ and $H^{(\sigma)}$ in Eqs. (49) and (50) are expressed in terms of the LECs given in Sect. II.1;
$G_{e}^{(\sigma)}$ $=\begin{dcases}-e_{1}+2e_{2}&(\sigma=1/2),\\ -e_{1}-e_{2}&(\sigma=3/2),\end{dcases}$ (62)
$G_{f}^{(\sigma)}$ $=\begin{dcases}\dfrac{-f_{1}-f_{2}-6f_{3}+2f_{4}+6f_{5}}{8}&(\sigma=1/2),\\ \dfrac{f_{1}+f_{2}-2f_{4}}{4}&(\sigma=3/2),\end{dcases}$ (63)
$H^{(\sigma)}$ $=\begin{dcases}2E_{1}-16E_{2}+E_{3}-3G_{1}&(\sigma=1/2),\\ 2E_{1}+8E_{2}+E_{3}-3G_{1}&(\sigma=3/2).\end{dcases}$ (64)

### v2

@@ Methods
[…]
@@ Conclusions
Since the correlation function shows limited sensitivity to the short-range 3BF, complementary approaches may be necessary.
@@ I Introduction
Three-baryon forces (3BFs), though not directly measurable and inherently scheme-dependent, are essential for understanding hypernuclear phenomena. One notable example is the inclusion of a $\Lambda NN$ 3BF, modeled as a two-pion exchange (TPE) plus phenomenological repulsive term, in calculations of hyperneutron matter [28]. More recently, lattice simulations of hyperneutron matter revealed the effect of $\Lambda NN$ and $\Lambda\Lambda N$ 3BFs, which are defined under the pionless effective-field theory (EFT) [42, 43].
@@
To investigate 3BFs with S=2S=-2, $\Xi NN$ systems serve as promising candidates. So far, various aspects of the $\Xi NN$ systems, such as bound states [11, 10, 9, 12, 6, 31, 5], resonant states [14, 13], and scattering [3], have been predicted based on the Faddeev approaches [30, 18]. In addition, the Gaussian expansion method [21] has been employed for studying the $\Xi NN$ bound states [22]. However, none of these studies take into account the $\Xi NN$ 3BF.
@@
This paper is organized as follows: In Sec. II, we formulate the general expression of the $\Xi NN$ 3BF as potentials in momentum space. We also give formulas necessary for implementing the potentials in the calculations of the correlation function. In Sec. III, we demonstrate how the $\Xi NN$ 3BF affects the correlation function. Section IV is devoted to the summary and perspectives of this work.
@@ II.1.2 Two-pion exchange term
$v_{\mathrm{TME}}$ $=-\dfrac{1}{4f_{0}^{4}}\dfrac{\left(\bm{\sigma}_{i}\cdot\bm{q}_{li}\right)\left(\bm{\sigma}_{k}\cdot\bm{q}_{nk}\right)}{\left(q_{li}^{2}+m_{\phi_{1}}^{2}\right)\left(q_{nk}^{2}+m_{\phi_{2}}^{2}\right)}\left[N_{1}^{\prime}+N_{2}^{\prime}\bm{q}_{li}\cdot\bm{q}_{nk}+N_{3}^{\prime}i\bm{\sigma}_{j}\cdot(\bm{q}_{li}\times\bm{q}_{nk})\right],$ (2)
@@
with the exchange operator $\mathcal{P}_{ij}=\mathcal{P}^{(q)}_{ij}\mathcal{P}^{(\sigma)}_{ij}\mathcal{P}^{(\tau)}_{ij}$}, where $\mathcal{P}^{(q)}_{ij}$, $\mathcal{P}^{(\sigma)}_{ij}$, and $\mathcal{P}^{(\tau)}_{ij}$, respectively exchange the parity, spin, and isospin of the subsystem formed by baryons $i$ and $j$. The pion-decay constant with the finite pion mass is expressed by $f_{\pi}$, for which we employ $f_{\pi}=92.4$ MeV, and $m_{\pi}=138.04$ MeV is the average pion mass. The axial-vector coupling constants $g_{A}$ and $g_{B}$ are expressed as $g_{A}=D+F$ and $g_{B}=-D+F$ with $D\approx 0.8$ and $F\approx 0.5$ [37]. The LECs $c_{i}$ and $u_{i}$ are given as a linear combination of other LECs;
@@ II.1.3 One-pion exchange plus contact term
$X_{123}^{456}$ $=\dfrac{g_{B}d}{2f_{\pi}^{2}}\dfrac{\bm{\sigma}_{1}\cdot\bm{q}_{41}}{q_{41}^{2}+m_{\pi}^{2}}\left[(\bm{\tau}_{2}-\bm{\tau}_{3})\cdot\bm{\tau}_{1}\,(\bm{\sigma}_{2}-\bm{\sigma}_{3})\cdot\bm{q}_{41}+(\bm{\tau}_{1}\times\bm{\tau}_{2})\cdot\bm{\tau}_{3}\,(\bm{\sigma}_{2}\times\bm{\sigma}_{3})\cdot\bm{q}_{41}\right],$ (16)
$X_{231}^{564}$ $=\dfrac{g_{A}}{2f_{\pi}^{2}}\dfrac{\bm{\sigma}_{2}\cdot\bm{q}_{52}}{q_{52}^{2}+m_{\pi}^{2}}\left[(e_{1}\bm{\sigma}_{3}+e_{2}\bm{\sigma}_{1})\cdot\bm{q}_{52}\,\bm{\tau}_{2}\cdot\bm{\tau}_{3}+(e_{3}\bm{\sigma}_{3}+e_{4}\bm{\sigma}_{1})\cdot\bm{q}_{52}\,\bm{\tau}_{1}\cdot\bm{\tau}_{2}-e_{5}(\bm{\tau}_{2}\times\bm{\tau}_{3})\cdot\bm{\tau}_{1}\,(\bm{\sigma}_{3}\times\bm{\sigma}_{1})\cdot\bm{q}_{52}\right],$ (17)
$X_{312}^{645}$ $=\dfrac{g_{A}}{2f_{\pi}^{2}}\dfrac{\bm{\sigma}_{3}\cdot\bm{q}_{63}}{q_{63}^{2}+m_{\pi}^{2}}\left[(e_{4}\bm{\sigma}_{1}+e_{3}\bm{\sigma}_{2})\cdot\bm{q}_{63}\,\bm{\tau}_{3}\cdot\bm{\tau}_{1}+(e_{2}\bm{\sigma}_{1}+e_{1}\bm{\sigma}_{2})\cdot\bm{q}_{63}\,\bm{\tau}_{2}\cdot\bm{\tau}_{3}-e_{5}(\bm{\tau}_{3}\times\bm{\tau}_{1})\cdot\bm{\tau}_{2}\,(\bm{\sigma}_{1}\times\bm{\sigma}_{2})\cdot\bm{q}_{63}\right],$ (18)
$X_{312}^{564}$ $=\dfrac{g_{A}}{2f_{\pi}^{2}}\dfrac{\bm{\sigma}_{3}\cdot\bm{q}_{63}}{q_{63}^{2}+m_{\pi}^{2}}\left[\{f_{1}\bm{\sigma}_{1}+f_{2}\bm{\sigma}_{2}+f_{3}i(\bm{\sigma}_{1}\times\bm{\sigma}_{2})\}\cdot\bm{q}_{63}\,\bm{\tau}_{3}\cdot\bm{\tau}_{1}+\{f_{2}\bm{\sigma}_{1}+f_{1}\bm{\sigma}_{2}+f_{3}i(\bm{\sigma}_{1}\times\bm{\sigma}_{2})\}\cdot\bm{q}_{63}\,\bm{\tau}_{2}\cdot\bm{\tau}_{3}\right$.
@@
$X_{231}^{645}$ $=\dfrac{g_{A}}{2f_{\pi}^{2}}\dfrac{\bm{\sigma}_{2}\cdot\bm{q}_{52}}{q_{52}^{2}+m_{\pi}^{2}}\left[\{f_{1}\bm{\sigma}_{3}+f_{2}\bm{\sigma}_{1}-f_{3}i(\bm{\sigma}_{3}\times\bm{\sigma}_{1})\}\cdot\bm{q}_{52}\,\bm{\tau}_{2}\cdot\bm{\tau}_{3}+\{f_{2}\bm{\sigma}_{3}+f_{1}\bm{\sigma}_{1}-f_{3}i(\bm{\sigma}_{3}\times\bm{\sigma}_{1})\}\cdot\bm{q}_{52}\,\bm{\tau}_{1}\cdot\bm{\tau}_{2}\right$.
@@ II.1.4 Contact term
$V_{ijk}^{a}=\begin{dcases}1&(a=1),\\ \bm{\sigma}_{j}\cdot\bm{\sigma}_{k}&(a=2)\\ \bm{\sigma}_{k}\cdot\bm{\sigma}_{i}&(a=3)\\ \bm{\sigma}_{i}\cdot\bm{\sigma}_{j}&(a=4)\\ i\bm{\sigma}_{i}\cdot\left(\bm{\sigma}_{j}\times\bm{\sigma}_{k}\right)&(a=5).\end{dcases}$ (34)
@@
$+\left[F_{1}-F_{2}(\bm{\sigma}_{2}-\bm{\sigma}_{3})\cdot\bm{\sigma}_{1}-F_{1}\bm{\sigma}_{2}\cdot\bm{\sigma}_{3}\right]\bm{\tau}_{1}\cdot\bm{\tau}_{3}+\left[G_{1}-E_{2}(\bm{\sigma}_{2}+\bm{\sigma}_{3})\cdot\bm{\sigma}_{1}-\dfrac{1}{3}E_{1}\bm{\sigma}_{2}\cdot\bm{\sigma}_{3}\right]\bm{\tau}_{2}\cdot\bm{\tau}_{3}$
$+F_{2}\left(\bm{\sigma}_{1}\times\bm{\sigma}_{2}\right)\cdot\bm{\sigma}_{3}\left(\bm{\tau}_{1}\times\bm{\tau}_{2}\right)\cdot\bm{\tau}_{3}.$ (35)
The LECs $E_{i}$, $F_{i}$, and $G_{1}$ can be expressed as a linear combination of the other LECs $C_{i}$, which enter the particle-basis Lagrangian [37]:
@@
$W^{(\sigma)}(\bm{r},\bm{R})$ $=\frac{1}{(2\pi)^{6}}\iiint d\bm{q}_{1}d\bm{q}_{2}d\bm{q}_{3}\exp\left[i(\bm{q}_{1}\cdot\bm{r}_{1}+\bm{q}_{2}\cdot\bm{r}_{2}+\bm{q}_{3}\cdot\bm{r}_{3})\right]\delta(\bm{q}_{1}+\bm{q}_{2}+\bm{q}_{3})$
$\times\Braket{\Upsilon_{\sigma m_{\sigma}}\Theta_{1/2,-1/2}|V_{(0)}^{\Xi NN}(\bm{q}_{1},\bm{q}_{2},\bm{q}_{3})|\Upsilon_{\sigma m_{\sigma}}\Theta_{1/2,-1/2}}.$ (46)
@@
$U_{\mathrm{TPE}}^{(u_{1})}(R)$ $=\frac{g_{A}^{2}u_{1}m_{\pi}^{6}}{32\pi^{2}f_{\pi}^{4}}\sum_{\lambda_{1}=0,1}\sum_{\lambda_{2}=0,1}\sum_{l}(-)^{\lambda_{1}}R^{\lambda_{1}+\lambda_{2}}\widehat{1-\lambda_{1}}\widehat{1-\lambda_{2}}\left[\left(\matrixquantity{3\\ 2\lambda_{1}}\right)\left(\matrixquantity{3\\ 2\lambda_{2}}\right)\right]^{\frac{1}{2}}$
$\times(1-\lambda_{1},0,1-\lambda_{2},0|l0)(\lambda_{1}0\lambda_{2}0|l0)\left\{\matrixquantity{1-\lambda_{1}&\lambda_{1}&1\\ \lambda_{2}&1-\lambda_{2}&l}\right\}$
@@
$U_{\mathrm{OPE}}^{(\sigma)}(R)$ $=-\dfrac{g_{A}G_{e}^{(\sigma)}m_{\pi}^{3}}{16\pi f_{\pi}^{2}}\int dr\varphi^{2}(r)\int_{-1}^{1}dw\left\{Y(m_{\pi}r_{2})f_{R_{0}}(r_{2})D_{R_{0}}(r_{3})+Y(m_{\pi}r_{3})f_{R_{0}}(r_{3})D_{R_{0}}(r_{2})\right\}$
$\quad-\dfrac{3g_{A}G_{f}^{(\sigma)}m_{\pi}^{3}}{8\pi f_{\pi}^{2}}\int dr\varphi^{2}(r)\int_{-1}^{1}dw\left\{Y(m_{\pi}r_{2})f_{R_{0}}(r_{2})D_{R_{0}}(r)+Y(m_{\pi}r_{3})f_{R_{0}}(r_{3})D_{R_{0}}(r)\right\},$ (49)
@@
Accordingly, the delta functions, which originate from the contact vertices [see Eqs. (105) and (106) in Appendix B], are regularized similarly,
@@
The newly defined LECs $G_{e}^{(\sigma)}$, $G_{f}^{(\sigma)}$ and $H^{(\sigma)}$ in Eqs. (49) and (50) are expressed in terms of the LECs given in Sec. II.1:
$G_{e}^{(\sigma)}$ $=\begin{dcases}-e_{1}+2e_{2}&(\sigma=1/2)\\ -e_{1}-e_{2}&(\sigma=3/2),\end{dcases}$ (62)
$G_{f}^{(\sigma)}$ $=\begin{dcases}\dfrac{-f_{1}-f_{2}-6f_{3}+2f_{4}+6f_{5}}{8}&(\sigma=1/2)\\ \dfrac{f_{1}+f_{2}-2f_{4}}{4}&(\sigma=3/2),\end{dcases}$ (63)
$H^{(\sigma)}$ $=\begin{dcases}2E_{1}-16E_{2}+E_{3}-3G_{1}&(\sigma=1/2)\\ 2E_{1}+8E_{2}+E_{3}-3G_{1}&(\sigma=3/2).\end{dcases}$ (64)